What is the point-slope form of the line that passes through (3,-3) and (5,1)?
Answer by CharlesG2(834)

What is the point-slope form of the line that passes through (3,-3) and (5,1)?

slope m = rise/run = (y2 - y1)/(x2 - x1)
m = (1 - -3)/(5 - 3) = (1 + 3)/2 = 4/2 = 2
point-slope form of a line:
y - y1 = m(x - x1) where (x1, y1) is the point to plug in, should be same result with both points given
using (5,1): y - 1 = 2(x - 5)
y - 1 = 2(x - 5)
y - 1 = 2x - 10
y = 2x - 9
using (3,-3): y - -3 = 2(x - 3)
y + 3 = 2(x - 3)
y + 3 = 2x - 6
y = 2x - 9, yes